### v2

# Phase space and data analyses of a nonminimally coupled scalar field system with a decaying dark energy model

Abstract
We investigate a class of scalar field dark energy models non-minimally coupled to gravity, characterized by a double exponential potential and parameterized coupling $\xi$. We study the cosmological dynamics for a recently proposed descending dark energy model, namely, Q-SC-CDM. Initially, we choose distinct values of coupling parameter. For some values of $\xi$, the evolution of the universe is split up into three different phases: decelerated expansion (early time), accelerated expansion (late-time) and slow-contraction (future era), and provide Big Crunch Singularity at distant future. In other scenario, the phase of slow-contraction vanishes, cosmic acceleration is obtained at current epoch, and the universe gets de-Sitter expansion at distant future. It is remarkable to see that the Big Crunch Singularity is redundant in the later case. Next, we investigate the phase space analysis for the model under consideration. Our investigation brings new asymptotic regimes and finds stable de-Sitter solution. Eventually, we perform a comprehensive Bayesian analysis using recent cosmological observations, including Cosmic Chronometers, Type Ia Supernovae (Pantheon+ and DES-SN5YR), and Baryon Acoustic Oscillation (DESI DR2) data. The results demonstrate that the present model yield constraints on key cosmological parameters—$\Omega_{0m}$, $H_{0}$ and the sound horizon $r_{d}$—that are consistent with $\Lambda$CDM within 68% confidence level, yet exhibit mild tension with Pantheon+ measurements. Additionally, we employ the $Om(z)$ diagnostic test, Akaike and Bayesian Information Criteria to distinguish our model from $\Lambda$CDM. The Statistical comparison reveals moderate support for the current model.
I Introduction
In recent years, scalar fields have become increasingly significant in cosmology and have an equation of state (EoS) parameter less than $-1/3$ to characterize the cosmic acceleration. They are used in various phenomenological models such as quintessence [1, 2, 3, 4]. In this case, a scalar field rolls down a potential energy function $V(\phi)$ that decreases monotonically over time. A wide range of different types of potential have been studied in the literature [5, 6, 7, 8, 9, 10, 11, 12]. In Planck mass units, the current value of $V(\phi)$ is minuscule. This indicates that the scalar field $\phi$ has a very low energy. There may be a potential that is currently shallow but will eventually reach more negative values. A specific configuration of $V(\phi)$ governs the dynamics of the scalar field and its influence on cosmic expansion. As the scalar field progresses, it can initiate various phases of the universe, ranging from accelerated expansion to gradual contraction. Recent studies have indicated that quintessence models featuring potentials that transition to negative values may lead to a future collapse of the universe, referred to as Big Crunch Singularity [13, 14, 15, 16, 17]. Alternatively, the universe may either collapse or enter a sequence of cycles characterized by an initial expansion followed by a contraction. Cyclic cosmology posits that the universe undergoes an infinite number of cycles, each commencing with a ‘Big Bang’ and concluding with a ‘Big Crunch’. Thus, the universe does not originate from a singularity; instead, it begins with a ‘Big Bang’ that leads to a temporary expansion before contraction ensues. The fundamental concept of the cyclic model is that the universe can circumvent issues associated with the initial singularity by experiencing an endless series of cycles, each of which eliminates any anomalies that arose in the previous cycle. For instance, in the early universe, the scalar field may not dominate the energy density, leading to conventional cosmological evolution, such as radiation dominance followed by matter dominance. However, as the scalar field progresses down its potential, it may eventually come to dominate, resulting in accelerated expansion. These transitions from matter dominance to accelerated expansion can occur in a smooth and continuous manner, governed by the evolution of the scalar field. This ensures that the evolution of the universe remains steady and gradual, devoid of sudden or disruptive changes. As the scalar field descends the potential, the kinetic energy rises while the potential energy diminishes. Eventually, the kinetic energy surpasses the potential energy. At this juncture, the total energy density of the universe reaches zero, indicating that the Hubble parameter, which quantifies the rate of expansion of the universe, also becomes zero. Consequently, the expansion of the universe ceases entirely and begins to contract. The contraction phase is anticipated to be gradual, implying that the universe will continue to contract at a diminishing rate over an extended duration.
In the standard quintessence scenario, the energy density of the scalar field minimally coupled to gravity mimics the effective cosmological constant. The detailed evolution is obviously dependent on a particular form of the potential, but the $\phi^{2}$ contribution can be regarded as a leading order term of expansion of the potential $V(\phi)$. We incorporate the non-minimal coupling constant to the quintessence scenario. Lykkas and Perivolaropoulos demonstrate that certain values of the non-minimal coupling parameter can prevent the cosmic apocalypse singularity in scalar-tensor quintessence with a linear potential [18]. For many years, researchers have examined the scalar tensor theories, which have a scalar field that is non-minimally coupled (NMC) to gravity. In order to reconcile Mach’s principle with general relativity, the first well-known instance of non-minimal coupling à $l$a Brans-Dicke theory was put forth in 1961 [19]. According to this theory, the gravitational constant is replaced by a scalar field $\phi$ entering the action in a particular combination with a Riemannian curvature as $\phi^{2}R$. Other types of scalar-tensor action were studied after the Brans-Dicke proposal: a well-known example of a NMC scalar field system is given by $F(\phi)R$ coupling with $F(\phi)=1-\xi\phi^{2}$. Such a theory has quite rich cosmic dynamics that need consideration. For a recent breakthrough, a NMC Higgs field because of a large coupling $\xi$ could result in a successful inflation, which is otherwise impossible [20]. The construction of a dark energy model is very interested in NMC scalar field systems because of their unique characteristics [21, 22, 23, 24, 25, 26, 27, 28, 29, 30, 31, 32, 33, 34, 35, 36, 37, 38, 39, 40, 41, 42, 43, 44, 45, 46, 47, 48, 49, 50, 51]. Non-minimal coupling, for example, may enable phantom crossover and result in cosmic scaling solutions that are relevant to dark energy models. The general characteristics of a NMC system with $F(\phi)=1-\xi\phi^{2}$ are phantom scaling solutions [52]. In the literature, two values of the coupling constant are frequently considered: $\xi=0$ (minimal coupling) and $\xi=1/6$ (conformal coupling). However, in many cases, $|\xi|\gg 1$ (strong coupling) has also been used [53, 54, 55, 56, 57, 58]. Recently, the value of $\xi$ was estimated from a distant supernova [59].
Dynamical system theory has been widely applied in cosmology to provide a broad understanding of dynamics for a variety of cosmological models. This approach has the benefit of having some sort of “machines” that uses a straightforward programmed algorithm to derive asymptotic solutions. In order to rewrite the original system as a system of first-order differential equations, a new set of variables must be introduced. This approach allows for the derivation of asymptotic solutions, and stability can be assessed using a straightforward programmed algorithm. Traditionally, physical cosmological solutions are derived from the combination of phase portraits and the stability of critical points. The ability to display any system trajectory that is permissible for every initial condition is the primary benefit of dynamical system analysis. Consequently, generic pathways to the accelerating phase (de-Sitter attractor) can be categorized. This paper aims to explore the cosmological dynamics of the Quintessence-Driven Slow-Contraction Cold Dark Matter (Q-SC-CDM) model and to examine its stationary points and stability. The Q-SC-CDM model was initially introduced by C. Andrei et al. in the framework of decaying dark energy and the end of cosmic expansion [60, 61]. In this research, we examine the cosmological dynamics of a NMC scalar field model with double exponential potential and a specific form of $F(\phi)$. We limit ourselves to the functional form $F(\phi)=1-\xi\phi^{2}$. We shall analyze the cosmological dynamics of the model under consideration and discuss the regions of deceleration, cosmic acceleration and slow-contraction. Furthermore, we shall conduct a phase space analysis of the Q-SC-CDM model, and investigate the existence of a stable late-time de-Sitter attractor solution that corresponds to an EoS of $-1$ and an energy density parameter $\Omega_{\phi}=1$. The paper is organized as follows. In section II, we discuss the equations of motion for a NMC scalar field system in a homogeneous and isotropic flat Friedmann–Lemaître–Robertson–Walker (FLRW) universe. The subsection II.1 is devoted to the cosmological dynamics of NMC system for different values of $\xi$, and explored the phases of deceleration, cosmic acceleration and contraction of the universe. In section III and subsection III.1, we construct the autonomous system that is useful for the dynamical system analysis, and discuss the stationary points, stability and cosmological importance for Q-SC-CDM model. In section IV, we reformulate the equations of motion to enhance the numerical stability. The sections V and VI are devoted to the data analysis and their results and discussion, respectively. The conclusion is presented in section VII.
II Equations of motion
A NMC scalar field model refers to a theoretical framework in which a scalar field is coupled to gravity not only through the usual minimal kinetic and potential terms, but also directly to the Ricci scalar $R$ of spacetime. This coupling modifies the Einstein field equations and allows for richer dynamics in the evolution of the universe. A common choice is $F(\phi)=1-\xi\phi^{2}$, where the coupling constant $\xi$ is a dimensionless parameter. The action for such a model typically takes the form [62, 63, 64, 65, 66, 67].
$S=\frac{1}{2}\int{\sqrt{-g}d^{4}x\Big{[}\frac{1}{\kappa^{2}}R-\Big{(}g^{\mu\nu}\phi_{\mu}\phi_{\nu}+\xi R\phi^{2}+2V(\phi)\Big{)}\Big{]}}+S_{M},$ (1)
where $S_{M}$ represents the action of matter, $\kappa^{2}=8\pi G=1/M_{Pl}^{2}$, $F(\phi)$ is a function of the scalar field representing the non-minimal coupling, $V(\phi)$ denotes the potential of the scalar field, $g$ is the determinant of the metric tensor $g_{\mu\nu}$. Varying the action (1) yields the equations of motion in a spatially flat FLRW background, which are given by
$H^{2}=\frac{\kappa^{2}}{3}\left(\frac{1}{2}{\dot{\phi}}^{2}+V(\phi)+6\xi H\phi\dot{\phi}+3\xi H^{2}\phi^{2}+\rho_{m}\right),$ (2)
$R=\kappa^{2}\left(-{\dot{\phi}}^{2}+4V(\phi)+18\xi H\phi\dot{\phi}+\xi R\phi^{2}+6\xi{\dot{\phi}}^{2}+6\xi\phi\ddot{\phi}+\rho_{m}(1-3\omega_{m})\right),$ (3)
$\frac{\ddot{a}}{a}$ $=$ $-\frac{\kappa^{2}}{3}\left((1-3\xi)\dot{\phi}^{2}-V(\phi)-3\xi H\phi\dot{\phi}-3\xi\phi\ddot{\phi}-3\xi\left(\frac{\ddot{a}}{a}\right)\phi^{2}+\frac{\rho_{0m}}{2a^{3}}\,\right),$ (4)
$\ddot{\phi}+3H\dot{\phi}+\xi R\phi+V^{\prime}(\phi)=0.$ (5)
Where Ricci scalar $R=6(2H^{2}+\dot{H})$. The pressure and energy density of matter are denoted by $p_{m}$ and $\rho_{m}$, and are related as $p_{m}=\omega_{m}\rho_{m}$; the $\omega_{m}$ represents the EoS for matter. The energy density and pressure for a NMC scalar field model are defined as
$\rho_{\phi}$ $=$ $\frac{1}{2}{\dot{\phi}}^{2}+V(\phi)+6\xi H\phi\dot{\phi}+3\xi H^{2}\phi^{2},$ (6)
$p_{\phi}$ $=$ $\frac{1}{2}{\dot{\phi}}^{2}-V(\phi)-\xi\left(4H\phi\dot{\phi}+2\dot{\phi}^{2}+2\phi\ddot{\phi}+(2\dot{H}+3H^{2})\phi^{2}\right).$ (7)
From equation (2), we get following constraint equation.
$\frac{\kappa^{2}\rho_{\phi}}{3H^{2}}+\frac{\kappa^{2}\rho_{m}}{3H^{2}}$ $=$ 11 11 (8)
$\Omega_{\phi}+\Omega_{m}$ $=$ 11 11 (9)
where $\Omega_{\phi}=\frac{\kappa^{2}\rho_{\phi}}{3H^{2}}$ and $\Omega_{m}=\frac{\kappa^{2}\rho_{m}}{3H^{2}}$ represent the energy density parameter for scalar field and matter, respectively. The EoS $w_{\phi}$ and total EoS $w_{eff}$ for a NMC scalar field model are given as
$w_{eff}$ $=$ $-1-\frac{2\dot{H}}{3H^{2}}$ (10)
$w_{\phi}$ $=$ $\frac{w_{eff}-w_{m}\Omega_{m}}{1-\Omega_{m}}$ (11)
Figure 1: The figure shows the evolution of potential (12) vs field. Initially, the potential is positive for negative values of scalar field. As field evolves, the universe expands with cosmic acceleration and remains so untill $\phi_{dec}$. On further evolution of field, the potential becomes more negative. To this effect, the expansion factor $a(t)$ collapses at $\phi_{con}$, and correspondingly Hubble parameter $H(t)$ becomes zero, see Fig. 2.
In the following subsection, we shall study the cosmological dynamics of decaying dark energy model with non-minimal coupling.
II.1 Cosmological Dynamics of Descending Dark Energy
The decaying dark energy model suggests that dark energy is not a constant but gradually loses energy over time. In such a model, the energy density of dark energy decreases as the universe evolves, either through a direct decay into other forms of energy or as a function of the scale factor of the universe. This is different from the standard cosmological model, which assumes that dark energy remains constant in its density (described by the cosmological constant $\Lambda$). Recently, a new model of decaying dark energy, namely, Q-SC-CDM was proposed by C. Andrei $et$ $al.$ [60]. The Q-SC-CDM scenario proposes that the universe underwent a period of slow-contraction rather than acceleration. In this case, quintessence acts as the driving force for the contraction and evolves in such a way that it causes the cosmic scale factor to shrink over time. We use the following potential proposed by C. Andrei $et$ $al.$ [60, 61].
$V(\phi)=V_{0}e^{-\kappa\alpha\phi}-V_{1}e^{\kappa\beta\phi},$ (12)
where $V_{0}$ and $V_{1}$ $(\alpha$ and $\beta)$ are constants of mass dimension four (zero). To study the cosmological behavior for NMC scalar field model, we use following dimensionless parameters
$H_{0}t$$\longrightarrow$ $t,$
$\kappa\phi$$\longrightarrow$ $\phi,$
$\frac{\kappa^{2}V_{0}}{H_{0}^{2}}$$\longrightarrow$ $V_{0}.$
$\frac{\kappa^{2}V_{1}}{H_{0}^{2}}$$\longrightarrow$ $V_{1}.$ (13)
The equations (4) and (5) in re-scaled form can be written as
$\frac{\ddot{a}}{a}=\frac{1}{1-\xi\phi^{2}}\left(-\frac{(1-3\xi)\dot{\phi}^{2}}{3}+\frac{V_{0}e^{-\alpha\phi}-V_{1}e^{\beta\phi}}{3}+\xi\frac{\dot{a}}{a}\phi\dot{\phi}+\xi\phi\ddot{\phi}-\frac{\Omega_{0m}}{2a^{3}}\right),$ (14)
$\ddot{\phi}+3\frac{\dot{a}}{a}\dot{\phi}+6\xi\left(\frac{\dot{a}^{2}}{a^{2}}+\frac{\ddot{a}}{a}\right)\phi-\alpha V_{0}e^{-\alpha\phi}-\beta V_{1}e^{\beta\phi}=0.$ (15)
Refer to caption Refer to caption Refer to caption
Refer to caption Refer to caption Refer to caption
[…]
Refer to caption Refer to caption Refer to caption
Refer to caption Refer to caption Refer to caption
Figure 3: The figure is similar to Fig. 2 except $\xi=0.5$. In this case, we get $q=-0.53$, $w_{eff}=-0.72$, $w_{\phi}=-0.98$ and $\Omega_{\phi}=0.70$ at current epoch. Moreover, the phase of slow-contraction vanishes, and we obtain cosmic acceleration at current epoch and de-Sitter expansion in distant future.
At early times, the universe was matter dominated and the field $\phi$ was frozen due to large Hubble damping. Therefore, it is straight forward to solve equations (14) and (15) numerically with the following initial conditions $(t\rightarrow t_{i}\simeq 0)$:
$a(t_{i})$ $=$ $\left(\frac{9\Omega_{0m}}{4}\right)^{1/3}t_{i}^{2/3},$
$\dot{a}(t_{i})$ $=$ $\frac{2}{3}\left(\frac{9\Omega_{0m}}{4}\right)^{1/3}t_{i}^{-1/3},$
$\phi(t_{i})$ $=$ $\phi_{i},$
$\dot{\phi}(t_{i})$ $=$ $0$ (16)
We consider a prior value of $\Omega_{0m}=0.3$. By tuning $\phi_{i}$, we get following parameters at the present epoch ($t_{0}$).
$a(t_{0})$ $=$ $1,$
$H(t_{0})$ $=$ $1$ (17)
where, $t_{0}$ is present time at which scale factor is unity. In Fig. 1, we exhibit the evolution of the potential (12) versus field $\phi$. Initially, potential shows positive behavior for negative scalar field, as time passes, the universe gets expansion with cosmic acceleration, and continues to evolve until $\phi_{dec}$ at which accelerated expansion turns decelerated expansion, on further evolution of $\phi$, the potential becomes more negative, and the scale factor $a(t)$ collapses ($t=t_{con}$) correspondingly Hubble parameter $H(t)$ goes to zero, see upper panels of Fig. 2. This implies that the expansion phase ends and slow-contraction begins. The lower panels of Fig. 2 display the evolution of the deceleration parameter $q(t)$, EoS $w(t)$ and energy density parameter $\Omega(t)$ versus cosmic time. Under the chosen model parameters with $\xi=0.1$, we get $q=-0.33$, $w_{eff}=-0.56$ and $w_{\phi}=-0.78$ at the present epoch. For accelerated expansion $w_{eff}$ should be less than $-1/3$. Though, we obtain an accelerated expansion, but the value of $w_{\phi}$ is not consistent with the present data of observation as according to the Planck 2018 results, $w_{\phi}=-1.00\pm 0.03$ [68]. However, for $\xi=0.5$ (rest of the parameters are same as mentioned in the caption of Fig. 2), we find $q=-0.53$, $w_{eff}=-0.72$, $w_{\phi}=-0.98$ and $\Omega_{\phi}=0.70$ at current epoch that are in good agreement with the present data [68], and the numerical evolution of cosmological parameters are shown in Fig. 3.
Let us make some comments on the value of non-minimal coupling $\xi$. As $\xi\rightarrow 0$, the standard quintessence scenario is recovered. In this case, the evolution of the universe is divided into three different phases such as deceleration (past era), acceleration (current epoch) and slow-contraction (future era) [60]. In quintessence models where potentials become negative may give rise to collapse of the universe in future, which is known as Big Crunch Singularity. Such singularity can be avoided in scalar-tensor theories with non-minimal coupling [18]. We numerically check the range of $\xi$ to alleviate the Big Crunch Singularity. For $0.20\leq\xi<$ 0.2, this kind of singularity can not be avoided whereas it vanishes for $\xi\geq 0.2$. In what follows let us consider two distinct cases one by one. As we move a little bit higher value of $\xi$ (e.g. $\xi=0.1$, see Fig. 2), we still obtain three different phases as mentioned earlier. However, the regime of slow-contraction is shifted in far future in comparison to the case of $\xi=0$ (standard quintessence). In other words, the Big Crunch Singularity is delayed for $\xi=0.1$. In case of $\xi=0.5$, the slow-contraction regime does not exist any more, an accelerated expansion is obtained for longer period, and finally the universe gets de-Sitter expansion in distant future, see Fig. 3. Therefore, the Big Crunch Singularity is redundant for large values of non-minimal coupling (e.g. $\xi=0.5$) in NMC scalar field model. The numerical evolution of scale factor $a(t)$, Hubble parameter $H(t)$ and deceleration parameter $q(t)$ of NMC scalar field model are well approximated by $\Lambda$CDM as shown in Fig. 3, where we get an accelerated expansion at current epoch and de-Sitter expansion in far future.
III Phase Space Analysis: Stationary points and their stability
Phase space analysis serves as a mathematical tool for examining the dynamics of systems by assessing the stability of critical points via eigenvalues and illustrating the trajectories within the phase space. It aids in comprehending how cosmological models evolve and remain stable. The fixed point can be classified as stable or unstable, depending on whether all eigenvalues are positive or negative. The eigenvalues at a certain fixed point can occasionally show both positive and negative signs. We call this location a saddle point. We present the subsequent dimensionless variables to facilitate the analysis of the dynamical system.
$x=\frac{\kappa\dot{\phi}}{\sqrt{6}H},\qquad y=\frac{\kappa\sqrt{V_{0}e^{-\kappa\alpha\phi}}}{\sqrt{3}H},\qquad A=\frac{\kappa\sqrt{V_{1}e^{\kappa\beta\phi}}}{\sqrt{3}H},\qquad z=\frac{\kappa\phi}{\sqrt{6}},$ (18)
The equations of motion (2)--(5) are reformulated in new variables as a system of differential equations of first order.
$\frac{dx}{dN}$ $=$ $x\Big{(}\frac{\ddot{\phi}}{H\dot{\phi}}-\frac{\dot{H}}{H^{2}}\Big{)},$
$\frac{dy}{dN}$ $=$ $-y\Big{(}\frac{\sqrt{6}x\alpha}{2}+\frac{\dot{H}}{H^{2}}\Big{)},$
$\frac{dA}{dN}$ $=$ $A\Big{(}\frac{\sqrt{6}x\beta}{2}-\frac{\dot{H}}{H^{2}}\Big{)},$
$\frac{dz}{dN}$ $=$ $x,$ (19)
where
$N$ $=$ $\ln a,$ (20)
$\frac{\dot{H}}{H^{2}}$ $=$ $-\frac{4+4A^{2}+2x^{2}-4y^{2}-12x^{2}\xi-24z^{2}\xi+6\sqrt{6}y^{2}z\alpha\xi+6\sqrt{6}A^{2}z\beta\xi+144z^{2}\xi^{2}-\Omega_{m}+3w_{m}\Omega_{m}}{2\left(1-6z^{2}\xi+36z^{2}\xi^{2}\right)},$
$\frac{\ddot{\phi}}{H\dot{\phi}}$ $=$ $-\frac{1}{2x\left(1-6z^{2}\xi+36z^{2}\xi^{2}\right)}\bigg{(}6x+\sqrt{6}y^{2}\alpha+\sqrt{6}A^{2}\beta-24A^{2}z\xi-12x^{2}z\xi+24y^{2}z\xi-36xz^{2}\xi$ (22)
$-6\sqrt{6}y^{2}z^{2}\alpha\xi-6\sqrt{6}A^{2}z^{2}\beta\xi+72x^{2}z\xi^{2}+216xz^{2}\xi^{2}+6z\xi\Omega_{m}-18w_{m}z\xi\Omega_{m}\bigg{)},$
$\Omega_{m}$ $=$ $1-(x^{2}+y^{2}-A^{2}+12\xi xz+6\xi z^{2}).$ (23)
We employ the autonomous system (19) to determine the stationary points by equating the left-hand side of these equations to zero. Consequently, we identify the following stationary points.
$A_{1}$.
$x=0,\qquad y=0,\qquad A=0,\qquad z=0.$ (24)
The corresponding eigenvalues are given by
${\mu}_{1}$ $=$ $3(1+w_{m})/2,$
${\mu}_{2}$ $=$ $3(1+w_{m})/2,$
${\mu}_{3}$ $=$ $-(3-3w_{m}+\sqrt{9-18w_{m}+9w_{m}^{2}-48\xi+144\xi w_{m}}),$
${\mu}_{4}$ $=$ $-(3-3w_{m}-\sqrt{9-18w_{m}+9w_{m}^{2}-48\xi+144\xi w_{m}}).$ (25)
Two eigenvalues are negative and the other two are positive. Therefore, this is a saddle point.
$A_{2}$.
$x=0,\qquad y=0,\qquad A=0,\qquad z=\pm\frac{1}{\sqrt{6}}.$ (26)
In this case, the eigenvalues are
${\mu}_{1}$ $=$ $2,$
${\mu}_{2}$ $=$ $2,$
${\mu}_{3}$ $=$ $-1,$
${\mu}_{4}$ $=$ $1-3w_{m}.$ (27)
This point is also saddle as one of the eigenvalue is negative and rests are positive.
$A_{3}$.
$x=0,\qquad y=\pm\frac{2\sqrt{-2\xi+\sqrt{\xi(\alpha^{2}+4\xi)}}}{\alpha},\qquad A=0,\qquad z=\frac{2\sqrt{6}\xi\pm\sqrt{6\alpha^{2}\xi+24\xi^{2}}}{6\alpha\xi}.$ (28)
The corresponding eigenvalues for $\alpha=\beta=\xi=1$ and $w_{m}=0$ are given by,
${\mu}_{1}$ $=$ $-3.316,$
${\mu}_{2}$ $=$ $-1.341+2.676i,$
${\mu}_{3}$ $=$ $-1.341-2.676i,$
${\mu}_{4}$ $=$ $-1.736\times 10^{-16}.$ (29)
All the eigenvalues exhibit negative behavior. Hence, it is a stable node. In this case, $\frac{\dot{H}}{H^{2}}=0$, $\Omega_{\phi}=1$ and $w_{eff}=w_{\phi}=-1$. Since,
$\frac{\dot{H}}{H^{2}}=0,$ (30)
We have
$a(t)$ $=$ $a_{0}e^{H_{0}(t-t_{0})}.$ (31)
In order to derive the expression for $\phi(t)$, we will utilize the subsequent dimensionless variable.
$x=\frac{\kappa\dot{\phi}}{\sqrt{6}H}$ (32)
At the stationary point where $x=0$, it follows that $\dot{\phi}=0$, leading to the conclusion that $\phi=\phi_{0}$ (constant). It is evident that both $\dot{H}=\dot{\phi}=0$, indicating that this point exhibits de-Sitter characteristics.
$A_{4}$.
$x=0,\qquad y=0,\qquad A=\pm\frac{2\sqrt{2\xi+\sqrt{\xi(\beta^{2}+4\xi)}}}{\beta},\qquad z=\frac{-2\sqrt{6}\xi\pm\sqrt{6\beta^{2}\xi+24\xi^{2}}}{6\beta\xi}.$ (33)
The corresponding eigenvalues for $\alpha=\beta=\xi=1$ and $w_{m}=0$ are given by,
${\mu}_{1}$ $=$ $-4.939,$
${\mu}_{2}$ $=$ $-0.530+1.113i,$
${\mu}_{3}$ $=$ $-0.530-1.113i,$
${\mu}_{4}$ $=$ $-3.132\times 10^{-16}.$ (34)
All eigenvalues display negative characteristics, indicating that it is a stable node. In this case, $\frac{\dot{H}}{H^{2}}=0$, $\Omega_{\phi}=1$ and $w_{eff}=w_{\phi}=-1$. Since,
$\frac{\dot{H}}{H^{2}}=0,$ (35)
We have
$a(t)$ $=$ $a_{0}e^{H_{0}(t-t_{0})}.$ (36)
Similar to point $A_{3}$, the point $A_{4}$ is also exhibited the de-Sitter behavior.
Refer to caption Refer to caption
Figure 4: The figure represents the stable fixed points $A_{3}$ (left panel) and $A_{4}$ (right panel) for $\alpha=\beta=\xi=1$ and $w_{m}=0$. The black dots represent stable attractor points.
III.1 Cosmological Importance
The critical points described above provide a roadmap for understanding the possible scenarios for the evolution of the universe. By analyzing their stability and cosmological implications, we can determine whether the universe is currently in an accelerating phase, a decelerating phase, or moving towards a stable attractor. Below, we shall break down the points in a more detailed manner, expanding on the differences in the cases and referring to the literature for comparison.
Point $A_{1}$ (Saddle Point): The fixed point $A_{1}$ behaves as a saddle, which means it has a combination of stable and unstable directions. The point corresponds to a non-accelerating universe with $\Omega_{\phi}=0$, indicating that the scalar field is not contributing to the energy density at this point, and the total EoS $w_{eff}=0$ signifies a universe that is neither accelerating nor decelerating in this regime. The saddle points often correspond to intermediate phases of the universe with energy components like matter dominating.
Point $A_{2}$ (Saddle Point): This point also exhibits saddle behavior with two eigenvalues positive and the other two negative. This suggests that it is unstable in one direction while stable in others. This point corresponds to accelerating or non-accelerating phase which depends on the value of $\xi$: $\Omega_{\phi}=\xi$ and $w_{eff}=-1+(1-\xi+8\xi^{2})/(1-\xi+6\xi^{2})$. The saddle point is often seen in models where different energy components (e.g., matter, radiation and dark energy) compete for dominance. The specific behavior of such point could vary based on the specific dynamics of the scalar field and matter content, but in general, saddle point like this is indicative of intermediate states in the evolution of universe.
Points $A_{3}$ and $A_{4}$ (Stable Attractor Solution): The fixed points $A_{3}$ and $A_{4}$ are stable attractors, which behave as a stable node with eigenvalues given by equations (III) and (III), respectively. It further shows that the solutions exhibit damped oscillatory behavior along the real parts of the eigenvalues with a stable attractior to the fixed point. At these points, the universe shows accelerating behavior with $\Omega_{\phi}=1$, meaning the scalar field dominates the energy density, while $\Omega_{m}=0$ indicates no matter content. The EoS is $w_{eff}=w_{\phi}=-1$, corresponding to a cosmological constant-like behavior, which is consistent with an accelerated expansion of the universe. The phase portraits provide crucial insight into the attractor behavior for both the points. As shown in Fig. 4, all trajectories approach the attractor point, which validates the stability of the fixed point and its role in describing the late-time behavior of the universe, particularly in the context of dark energy dominance. The stable node is typical of models where dark energy dominates at late-time. It corresponds to a cosmological constant or quintessence field behavior, where the universe enters a phase of accelerated expansion.
IV Equations for Data analysis
In this section, we reconstruct the first-order dynamical equations that facilitate the evaluation of the Hubble parameter in our model. In the previous section, we analyzed the qualitative behavior of the system using the dynamical systems framework. However, the equations expressed in that form exhibit numerical instability. Such instability is common in highly non-linear models; therefore, reformulating the system in a slightly different manner can significantly enhance numerical stability. It is important to note that the autonomous system presented earlier can still be used to derive constraints on the model parameters. In the previous works, the parameters for the $k-$essence and quintessence scalar field model have been constrained using the dynamical system analysis framework [69, 70, 71]. The reformulation presented here does not alter the physical interpretation or the parameter constraints of the system. Therefore, the redefined variables yield equivalent information. The redefined dimensionless dynamical variables are
$x=\frac{\kappa\dot{\phi}}{\sqrt{6}H_{0}},\quad\tilde{y}=\frac{\kappa\sqrt{V}}{\sqrt{3}H_{0}},\quad\tilde{\Omega}_{m}=\frac{\kappa\rho_{m}}{3H_{0}^{2}},\quad h=\frac{H}{H_{0}},\quad z=\kappa\phi,$ (37)
where the field and fluid variables are normalized using the Hubble constant $H_{0}$. In terms of these variables, the Friedmann equation becomes:
$h^{2}=x^{2}+\tilde{y}^{2}+2\sqrt{6}\xi xzh+\xi h^{2}z^{2}+\tilde{\Omega}_{m}\ .$ (38)
Solving the quadratic equation for $h$, we obtain the following valid root:
$h=-\frac{\frac{1}{2}\sqrt{24\xi^{2}x^{2}z^{2}-4\left(\xi z^{2}-1\right)\left(-A^{2}+\tilde{\Omega}_{m}+x^{2}+y^{2}\right)}+\sqrt{6}\xi xz}{\xi z^{2}-1}\ ,$ (39)
which ensures a positive value of $h$ throughout the evolution. Since $h$ is a monotonically increasing positive function of redshift, we discard the other root, which yields unphysical negative values of $h$. In this study, we adopt a dual-exponential form of the potential, characterized by two parameters:
$V(\phi)=V_{0}e^{-\kappa\alpha\phi}-V_{1}e^{\kappa\beta\phi}\ ,$ (40)
which results in $\tilde{y}$
$\tilde{y}^{2}=y^{2}-A^{2}\ ,$ (41)
where
$y^{2}=\frac{\kappa^{2}V_{0}e^{-\kappa\alpha\phi}}{3H_{0}^{2}},\quad A^{2}=\frac{\kappa^{2}V_{1}e^{\kappa\beta\phi}}{3H_{0}^{2}}\ .$ (42)
With this dynamical variables, the Ricci scalar can be expressed as:
$R=\frac{H_{0}^{2}}{1-\xi z^{2}}\left(-6x^{2}+12\tilde{y}^{2}+18\sqrt{6}\xi hzx+36\xi x^{2}+6\xi z\frac{\kappa\ddot{\phi}}{H_{0}^{2}}+3\tilde{\Omega}_{m}(1-3w_{m})\right)\ .$ (43)
For a flat FLRW metric, the Ricci scalar takes the form
$R=12H^{2}+6\dot{H}\ ,$ (44)
which allows us to express the time derivative of the Hubble parameter $\dot{H}$ by using the field equations, and given as:
$\dfrac{\dot{H}}{H^{2}}=-2+\frac{1}{6h^{2}(1-\xi z^{2})\left(1+\dfrac{6z^{2}\xi^{2}}{1-z^{2}\xi}\right)}\left(-6x^{2}+12\tilde{y}^{2}+36\xi x^{2}+18\xi z\alpha(A^{2}+y^{2})+3\tilde{\Omega}_{m}(1-w_{m})\right)\ .$ (45)
Analogous to the procedure in the preceding section, we derive the first-order autonomous system of equations from the field equations as:
$x^{\prime}$ $=$ $\frac{1}{\sqrt{6}}\left(\frac{1}{1+\frac{6\xi^{2}z^{2}}{1-\xi z^{2}}}\right)\bigg{(}-3\sqrt{6}x+\frac{3y^{2}\alpha}{h}+\frac{3A^{2}\beta}{h}-$ (46)
$\frac{\xi z}{h(1-\xi z^{2})}\left[-6x^{2}+12\tilde{y}^{2}+18\sqrt{6}z\xi hx+36\xi x^{2}+3\tilde{\Omega}_{m}(1-3\tilde{\Omega}_{m})\right]\bigg{)},$
$y^{\prime}$ $=$ $\frac{-1}{2h}y\sqrt{6}x\alpha,$ (47)
$A^{\prime}$ $=$ $\frac{1}{2h}A\beta\sqrt{6}x,$ (48)
$z^{\prime}$ $=$ $\frac{x\sqrt{6}}{h},$ (49)
$\tilde{\Omega}_{m}^{\prime}$ $=$ $-3\tilde{\Omega}_{m}(1+w_{m})\,,$ (50)
where $()^{\prime}=d()/dN$, $N$ representing the number of e-folds $N=\log a$. Unlike the previous case, where the dynamics of the matter energy density $\Omega_{m}$ were obtained using the Friedmann constraint Eq. (23), here we express the Hubble parameter $h$ in terms of the remaining dynamical variables. This reformulation increases the dimensionality of the phase space from 4 to 5. Since the primary objective in this section is to accurately capture the Hubble evolution and compare it with observational data, the extended phase space enables a more comprehensive exploration of the parameter space.
As we evolve the system toward higher redshifts, i.e., using the variable $N=-\ln(1+z)$, [^1] Here, $z$ denotes the redshift. This should not be confused with the dynamical variable $z=\kappa\phi$. From this point onward, whenever the term redshift is used, it will be denoted by $z$. we set initial conditions at the current epoch, corresponding to $N=0$. It immediately follows from the Hubble Eq. (38) that the condition $h(0)=1$ must be satisfied. As a result, the initial value of one of the dynamical variables—say $y$—can be expressed in terms of the others as:
$y(0)=\sqrt{1-x_{0}^{2}+A_{0}^{2}-2\sqrt{6}\xi x_{0}z_{0}-\xi z_{0}^{2}-\tilde{\Omega}_{0m}}\ .$ (51)
Thus, although the system consists of five equations, the Hubble constraint Eq. (38) effectively reduces the parameter space by one degree of freedom. This is a key equivalence between the current formulation and the previous one. In the next section, we will iterate over the remaining parameters to obtain the best-fit values. Additionally, we define the cosmological parameters, such as the fractional energy densities as follows:
$\Omega_{\phi}=\frac{\kappa^{2}\rho_{\phi}}{3H^{2}}=\frac{x^{2}+\tilde{y}^{2}}{h^{2}},\quad\Omega_{\rm m}=\frac{\kappa^{2}\rho_{m}}{3H^{2}}=\frac{\tilde{\Omega}_{m}}{h^{2}},\quad\Omega_{\rm eff}=\Omega_{\phi}+\dfrac{2\sqrt{6}\xi xzh+\xi h^{2}z^{2}}{h^{2}}=1-\Omega_{m}\ .$ (52)
Here, we adopt the minimally coupled analogy for the scalar field, where the energy density takes the standard form $\rho_{\phi}=\frac{1}{2}\dot{\phi}^{2}+V(\phi)$. To account for non-minimal coupling effects, we introduce an additional quantity; the effective energy density $\Omega_{\rm eff}$. Similarly, we define the equation of state of the scalar field as $\omega_{\phi}=p_{\phi}/\rho_{\phi}$, and the effective equation of state as:
$\omega_{\phi}=\frac{x^{2}-\tilde{y}^{2}}{x^{2}+\tilde{y}^{2}}\ ,\quad\omega_{\rm eff}=-1-\frac{2\dot{H}}{3H^{2}}\ ,$ (53)
where $\frac{\dot{H}}{H^{2}}$ is given by Eq. (45).
V Data Analysis
In this section, we discuss the datasets used to constrain the model parameters.
  • CC Data: The Cosmic Chronometer (CC) data set contains $32$ model independent observational data points of the Hubble parameter in the redshift range $\in[0.07,1.965]$ [72, 73, 74].
  • SN Data: In this study, we use observational data from Type Ia Supernovae (SNe) provided by two different observations: Pantheon Plus and DES5YR [75, 76, 77]. The most recent update from the Pantheon+ compilation includes $1701$ cosmologically viable SNe Ia light curves from $1550$ distinct supernovae. We refer to this as the “PP” sample, covering the redshift range $\in[0.001,2.26]$. We use the likelihood provided in Ref. [75], which incorporates the full statistical and systematic covariance matrix. From this catalog, we use the data column corresponding to the apparent magnitude of the supernovae $m$, rather than the distance modulus $\mu$.
    The second supernova dataset is from the Dark Energy Survey (DES-SN5YR), which includes $1829$ distinct SNe [77]. It consists of $194$ nearby SNe samples with redshift $<0.1$, and $1635$ DES SNe samples. We compute the likelihood using the data column corresponding to the distance modulus $\mu$, along with the full covariance matrix. [^2] The likelihood estimation and construction of covariance matrix is shown in https://github.com/des-science/DES-SN5YR/blob/main/5_COSMOLOGY/SN_only_cosmosis_likelihood.py Throughout the text, we refer to this dataset as “DES”. The distance modulus is defined as:
    $\mu=m-M=5\log(d_{L}/\text{Mpc})+25\ \ .$ (54)
    Here, $m$ is the apparent magnitude of the supernova and $d_{L}$ is the luminosity distance:
    $d_{L}({z})=c(1+{z})\int_{0}^{{z}}\frac{dz^{\prime}}{H(z^{\prime})}\ ,$ (55)
    for the flat FLRW metric, where $c$ denotes the speed of light in km/s. The model parameters are constrained by minimizing a chi-square ($\chi^{2}$) likelihood, defined as:
    $-2\ln(\mathcal{L})=\chi^{2}=\Delta D^{T}\mathcal{C}^{-1}_{\rm stat+sys}\Delta D\ ,$ (56)
    where
    $\Delta D=\mu_{\rm Obs}-\mu_{\rm Model}\ .$ (57)
  • BAO data: The Baryonic Acoustic Oscillation (BAO) dataset, referred to as DESBAO, consists of 7 data points from the Dark Energy Spectroscopic Instrument (DESI) Release II [78], which is an updated and enhanced version of DR1 [77, 80, 81]. We derive the constraints using the full covariance matrix provided in [78]. The relevant observables include $(D_{M},D_{H},D_{V},r_{d})$, corresponding to the comoving angular distance, comoving Hubble distance, spherically averaged distance, and the sound horizon at the drag epoch, respectively. For mathematical definitions of these parameters, we refer readers to Refs. [79, 77]. In this study, we treat $r_{d}$ as a free parameter, which we constrain using various combinations of datasets. Specifically, we consider two distinct combinations:
    $(i)\ \chi^{2}_{\rm tot}=\chi^{2}_{\rm CC}+\chi_{\rm DESBAO}^{2}+\chi^{2}_{\rm PP}\ ,\quad(ii)\ \chi^{2}_{\rm tot}=\chi^{2}_{\rm CC}+\chi_{\rm DESBAO}^{2}+\chi^{2}_{\rm DES}\ .$ (58)
    The value of $r_{d}$ as estimated from Planck observations using $\Lambda$CDM is $r_{d}=147.09\pm 0.26$ [68] where as $H_{0}r_{d}$ from DESI DR2 is approximately $(100-101)\times 100\ \mathrm{km\,s^{-1}}$, reflecting an approximate 40% improvement over DR1. The total likelihood of the system is then computed as:
    $-2\ln\mathcal{L}_{\rm tot}=\chi^{2}_{\rm tot}\ ,$ (59)
    from which we obtain the 1D and 2D posterior probability distributions of the model parameters by marginalizing over the absolute magnitude $M_{b}$ from the PP dataset. The posterior distribution is obtained using Bayes’ theorem, which requires careful selection of prior ranges. Based on our prior discussion of the system’s dynamics, we apply uniform priors as listed in Table 1. For sampling and likelihood estimation, we use the publicly available nested sampler PolyChord, which is well suited for high-dimensional parameter spaces compared to affine-invariant Markov Chain Monte Carlo (MCMC) ensemble samplers like emcee [82, 83, 84]. Our model is implemented in Cobaya (Code for Bayesian Analysis), an open-source Python framework for parameter inference, and the resulting samples are analyzed using GetDist [85, 86]. Finally, we compare the performance of the current model to flat $\Lambda$CDM using information criteria such as the Akaike Information Criterion (AIC) and the Bayesian Information Criterion (BIC) [87, 88, 89].
    $\rm{AIC}=-2\ln\mathcal{L}_{\rm max}+2k,$ (60)
    $\rm{BIC}=-2\ln\mathcal{L}_{\rm max}+k\ln N,$ (61)
    where $k$ is the number of model parameters, $N$ is the total number of data points, and $\mathcal{L}_{\rm max}$ is the maximum likelihood. To assess model preference, we compute the difference in information criteria (IC) between the current model and $\Lambda$CDM:
    $\Delta\rm IC=IC_{\rm Model}-IC_{\Lambda CDM}\ .$ (62)
    A model with a lower IC value is preferred. A difference of $\Delta\rm IC<2$ indicates substantial support for the model, values between $2$ and $7$ suggest weak support, and values $>10$ imply no support.
Parameters Prior range
$\Omega_{m}$ $[0,0.7]$
$H_{0}$ $[30.0,100.0]$
$\alpha$ $[-1.8,2.0]$
$\beta$ $[-1.6,4.0]$
$A_{0}$ $[0,3.0]$
$z_{0}$ $[-1.2,2.0]$
$r_{d}$ $[100.0,300.0]$
$M_{b}$ $[-20.0,-18.0]$
Table 1: The uniform prior range for models $\xi=0.1$ and 0.05.
VI Results and Discussion
The available cosmological data allow us to probe the model’s qualitative behavior up to the present epoch, but not into the distant future. In our earlier analysis, we considered two representative values of $\xi$—namely, 0.1 and 0.5—to investigate the avoidance of a Big Crunch singularity in the far future. Therefore, to ensure compatibility with observational data and avoid such redundancy, we focus on smaller values of $\xi$ when estimating the posterior distributions of the model parameters.
[…]
since the initial conditions are highly sensitive to the value of the non-minimal coupling parameter. Consequently, treating $\xi$ as a free parameter introduces stiffness into the differential equations, making it extremely challenging to constrain the remaining parameters reliably. We have also observed that to reproduce the late-time cosmic acceleration, the initial condition of the variable $x$ must be smaller than $0.1$. Therefore, we fix its initial value to $x(0)=10^{-4}$. The marginalized 1D and 2D posterior distributions of the model parameters for both combinations of datasets and both models are presented in Fig. 5, while the corresponding mean values at the $68\%$ confidence level are summarized in Table 2.
[…]
Parameter 68% limits 68% limits 68% limits 68% limits
$\Omega_{0m}$ $0.308^{+0.010}_{-0.0086}$ $0.3214\pm 0.0078$ $0.305^{+0.013}_{-0.0090}$ $0.319^{+0.010}_{-0.0078}$
$H_{0}$ $69.1\pm 1.7$ $69.57\pm 0.25$ $69.1\pm 1.8$ $69.57\pm 0.26$
$\alpha$ $-0.25^{+0.65}_{-0.41}$ $-0.27^{+0.65}_{-0.42}$ $-0.18^{+0.55}_{-0.35}$ $-0.23^{+0.57}_{-0.28}$
$\beta$ $0.9^{+1.2}_{-1.9}$ $1.0^{+1.5}_{-1.8}$ $0.96^{+1.2}_{-2.0}$ $0.99^{+1.3}_{-1.9}$
$A_{0}$ $0.64^{+0.20}_{-0.62}$ $0.71^{+0.15}_{-0.70}$ $0.68^{+0.12}_{-0.67}$ $0.54^{+0.14}_{-0.52}$
$z_{0}$ $0.03^{+0.37}_{-0.42}$ $0.02\pm 0.32$ $0.08^{+0.60}_{-0.78}$ $0.09^{+0.54}_{-0.63}$
$r_{d}$ $145.4^{+3.3}_{-3.8}$ $143.01^{+0.74}_{-0.83}$ $145.4^{+3.4}_{-4.0}$ $143.06\pm 0.78$
Table 2: The parameter estimation at $68\%$ confidence level.
Figure 5: The corner plot of decay dark energy model for $x_{0}=10^{-4}$.
We find that both models are consistent with the late-time observational datasets, yielding robust constraints on the matter energy density $\Omega_{0m}$, the Hubble parameter $H_{0}$ and the sound horizon scale $r_{d}$. These values are broadly consistent with the $\Lambda$CDM model and simultaneously exhibit a noticeable discrepancy with the SH0ES result [75]. In this study, we aim to assess the impact of supernova datasets from two distinct surveys, while keeping the BASE dataset (CC+DESBAO) fixed. We observe that the DES supernova data favor slightly higher values of $\Omega_{0m}$ for both models, while the inferred values of $H_{0}$ remain nearly unchanged. However, as $\Omega_{0m}$ increases, the inferred value of the sound horizon $r_{d}$ tends to decrease, indicating a mild tension—albeit at a small magnitude—between the two supernova datasets.
Figure 6: The correlation matrix for $\xi=0.1$ and $x_{0}=10^{-4}$ both the combined data set marginalizing $M_{b}$.
Additionally, we present the correlation matrix in Fig. 6, generated using Seaborn python package, utilizing the sample chains obtain from MCMC [90]. The matrix is constructed by combining samples from both datasets, as the parameter correlations are found to be qualitatively similar across them. In this matrix, higher positive (or negative) values indicate stronger correlations between the parameters. A positive correlation implies that as one parameter increases, the other tends to increase proportionally, whereas a negative correlation indicates inverse proportionality. For example, $H_{0}$ and $r_{d}$ exhibit a strong negative correlation, implying that an increase in $H_{0}$ corresponds to a decrease in $r_{d}$, and vice versa. A similar inverse correlation is observed between $\Omega_{0m}$ and $r_{d}$. In contrast, the correlation between $\Omega_{0m}$ and $H_{0}$ is weak and also negative. Among the model parameters, $\alpha$ and $\beta$, which define the structure of the potential, are strongly anti-correlated. Furthermore, $\alpha$ is positively correlated with $z_{0}$, the initial value of the field variable $z=\kappa\phi$, while showing negligible correlation with $\beta$. The parameter $A_{0}$, on the other hand, shows negative correlation with both $\alpha$ and $\beta$, suggesting a suppressive effect on the potential parameters.
Figure 7: The evolution of cosmological parameters corresponding to the best fit values obtained from MCMC.
The qualitative behavior of the models is illustrated in Fig. 7 as a function of redshift $z$, where we plot the fractional energy densities of the scalar field and matter fluid, the EoS of the field, and the effective EoS of the system, using the best-fit parameter values obtained from the MCMC analysis. The evolution of the field energy density indicates that it begins to dominate near the current epoch, with the corresponding EoS of the field approaching $\omega_{\phi}\approx-1$, while the effective EoS satisfies $\omega_{eff}<-0.5$, mimicking the behavior of the $\Lambda$CDM model. At present, the matter density is approximately $\Omega_{m}\sim 0.3$, and it increases at higher redshifts, reflecting the expected matter-dominated era where the effective EoS tends toward zero. As the system is extrapolated into the asymptotic past, the effective EoS rises toward $\omega_{eff}\to 1$, signaling a stiff matter-like behavior, with $\Omega_{eff}$ correspondingly dominating. It is important to note that, since radiation has not been included in this system, the onset of stiff matter behavior occurs at intermediate redshifts for specific parameter configurations. Inclusion of a radiation component would shift this stiff phase further back into the early universe. However, given that the observational data used in this study are largely limited to redshifts $z\lesssim 3$, the radiation density does not significantly affect our late-time cosmological analysis.
Figure 8: The $\mathcal{O}m(z)$ diagnosis of the current model for distinct data sets is plotted against the redshift $z$ and compared with $\Lambda$CDM.
We further evaluate an important diagnostic, known as the $Om(z)$, which is particularly effective in distinguishing dynamical dark energy models from the cosmological constant in the low-redshift regime. The $Om(z)$ function is defined as [91, 92, 93].
$Om(z)=\frac{h^{2}(z)-1}{(1+z)^{3}-1},\quad h(z)\equiv H(z)/H_{0}\ .$ (64)
For a spatially flat $\Lambda$CDM model, $Om(z)$ reduces to a constant equal to $\Omega_{0m}$. In contrast, deviations from constancy in $Om(z)$ signify the presence of dynamical dark energy. As shown in Fig. 8, the current model exhibits behavior very similar to that of $\Lambda$CDM within the redshift range $z<2$. Furthermore, for certain data set combinations (denoted D1 and D2), the models with fixed coupling parameters $\xi=0.1$ and $\xi=0.05$, respectively, continue to mimic $\Lambda$CDM even at higher redshifts, up to $z\sim 6$. This consistency across a broad redshift range supports the model’s viability in describing late-time cosmic acceleration.
Model AIC BIC $\Delta$AIC $\Delta$BIC $\chi_{\rm rd}^{2}$
$\xi=0.1$[D1] $1801.51$ $1845.20$ $5.82$ $22.20$ $1.03$
$\xi=0.1$[D2] $1712.82$ $1751.54$ $4.99$ $21.58$ $0.91$
$\xi=0.05$[D1] $1801.63$ $1845.33$ $5.94$ $22.33$ $1.03$
$\xi=0.05$[D2] $1712.12$ $1750.85$ $4.29$ $20.89$ $0.91$
$\Lambda$CDM [D1] 1795.69 1823.00 0 0 $1.03$
$\Lambda$CDM [D2] $1707.83$ $1729.96$ 0 0 $0.91$
Figure 9: The evaluation of AIC and BIC of the models are tabulated in the table and a graphical comparison of AIC and BIC with $\Lambda$CDM, where dashed lines is corresponding to $\Lambda$CDM, $\Delta$AIC/$\Delta$BIC =0 is shown in the right panel.
We assess the statistical performance of the proposed models using information criteria and compare them with the standard $\Lambda$CDM model as shown in Fig. 9. In addition, we compute the reduced chi-square statistic, defined as $\chi_{\rm rd}^{2}\equiv\chi_{\rm min}^{2}/\nu$, where $\nu$ represents the number of observational data points minus the number of fitted parameters. This metric provides an important measure of goodness-of-fit. A value of $\chi_{\rm rd}^{2}\ll 1$ typically indicates model overfitting, whereas $\chi_{\rm rd}^{2}\gg 1$ suggests a poor fit. A value near unity implies that the model fits the data well. Our results show that the proposed models yield $\chi_{\rm rd}^{2}\sim 1$, consistent with that of the $\Lambda$CDM model, indicating excellent agreement with the observational data. Furthermore, we find that the current models yield a $\Delta\rm AIC$ of approximately 5, suggesting moderate support for the models relative to $\Lambda$CDM, though not sufficient for a strong preference. Conversely, the $\Delta\rm BIC>10$ indicates weak support for the current models when penalizing model complexity more strongly. Among the two considered cases, the model with $\xi=0.1$ exhibits a slightly better performance in both AIC and BIC metrics compared to the model with $\xi=0.05$, thus indicating a mild statistical preference.
VII Conclusion
In this paper, we used Q-SC-CDM model to study the cosmological dynamics with non-minimal coupling. We started the detailed investigations of the dynamics with a well known example of $F(\phi)R$ coupling, where $F(\phi)=1-\xi\phi^{2}$. The numerical evolution of Q-SC-CDM potential was shown in Fig. 1. Initially, the potential $V(\phi)$ exhibited the positive behavior for negative values of scalar field. As field evolved, the universe obtained an accelerated expansion and remained so untill $\phi_{dec}$. The potential showed negative behavior on the further evolution of field that led to the collapse of the universe, referred to as Big Crunch Singularity. We numerically checked the range of $\xi$ to avoid such singularity. It was noticed that the Big Crunch Singularity remained present in the range of $0.20\leq\xi<$ 0.2 whereas it was completely vanished for $\xi\geq 0.2$. In our work, we chose two distinct values of $\xi$ such as 0.1 and 0.5. For $\xi\rightarrow 0$, NMC scalar field model was reduced to the standard quintessence and Big Crunch Singularity remained present. In case of $\xi=0.1$, we numerically evolved the evolution equations (14) and (15). The results were depicted in Fig. 2. The numerical evolution of the universe was divided into three distinct regions: decelerated expansion, accelerated expansion and slow-contraction. In this case, Big Crunch Singularity is delayed in distant future but not completely avoided, see Fig. 2. For $\xi=0.5$, the slow-contraction phase vanished. Consequently, the Big Crunch Singularity was redundant. Therefore, we obtained an accelerated expansion for a longer period and de-Sitter expansion in distant future. The numerical evolution of different cosmological parameters were shown in Fig. 3 which were well approximated by $\Lambda$CDM model.
Next, we constructed an autonomous system by using a convenient set of dimensionless variables to performed the phase space analysis. We obtained stationary points, discussed their stabilities by looking at the eigenvalues, and found asymptotic regimes of solutions in the model. The critical point $A_{1}$ behaved as a saddle, and corresponded to a non-accelerating universe with $\Omega_{\phi}=0$ and $w_{eff}=0$. As a result, the scalar field was not contributing to the energy density at this point. The fixed point $A_{2}$ was also exhibited saddle behavior with two eigenvalues positive and the other two negative. The fixed points $A_{3}$ and $A_{4}$ were behaved as a stable node, and exhibited accelerated expansion with $\Omega_{\phi}=1$ and $w_{eff}=w_{\phi}=-1$, which mimicked the cosmological constant-like behavior. The phase portraits for both points were shown in Fig. 4, where all trajectories approached the stable attractor points.
Finally, we constrained the model parameters by fixing the non-minimal coupling parameter $\xi$ to two representative values: $\xi=0.1$ and 0.05. These values were chosen to avoid numerical stiffness encountered while solving the equations. We also observed that increasing $\xi$ beyond 0.1 leads to divergence in other dynamical variable, making the numerical evolution extremely difficult. Therefore, our comparison is restricted to these two cases. The models were tested using two distinct supernova datasets—Pantheon+ and DES—while keeping the base datasets (CC + DESBAO) fixed. The results indicate that both models exhibit similar behavior when applied to the same dataset and show a notable departure from the standard $\Lambda$CDM model at higher redshifts. This deviation is characterized using the $Om(z)$ diagnostic, which reveals that for the PP dataset, the effective fluid energy density in both models resembles that of a cold dark matter component. However, for the DES dataset, the deviation becomes significantly more pronounced. Additionally, we assessed the statistical significance of the models relative to $\Lambda$CDM using the AIC and BIC. The AIC results indicate moderate support for both models, whereas the BIC suggests only weak support in comparison to the fiducial model.
Acknowledgments
This research was funded by the Science Committee of the Ministry of Science and Higher Education of the Republic of Kazakhstan (Grant No. AP22682760). SH acknowledges the support of National Natural Science Foundation of China under Grants No. W2433018 and No. 11675143, and the National Key Research and development Program of China under Grant No. 2020YFC2201503. MS acknowledges Integral University, Lucknow for financial support through Seed Money Grant 2024-2025 (Project Sanction No.: IUL/ICEIR/SMP/2024-04) and MCN: IU/R&D/2025-MCN0003753. MS also thanks the Inter-University Centre for Astronomy and Astrophysics (IUCAA), Pune for the hospitality and facilities under the visiting associateship program where the work was completed.